OpenStudy (anonymous):

Prove that the two tangents to a parabola from any point on the directrix are perpendicular.

OpenStudy (anonymous):

To put it into context, I know that..\[\frac{ d }{ dx }x^2=2x\]..and that the directrix of that parabola is..\[(0,-\frac{ 1 }{ 4 })\]Please help me with this proof...

OpenStudy (anonymous):

The statement doesn't make sense to me...

OpenStudy (anonymous):



OpenStudy (anonymous):